A car travels one-third of the path with the speed of 40 km/h, one-fourth of the path with a speed of 30 km/h and the remaining part of the path with a speed of 50 Km/h. What is the avarege speed of the car for the whole trip in km/h


1
Expert's answer
2021-05-25T10:25:54-0400

"v_{avr}=\frac{s/3+s/4+5s/12}{s/(3\cdot40)+s/(4\cdot30)+5s/(12\cdot50)}="


"=\frac{1}{1/(3\cdot40)+1/(4\cdot30)+5/(12\cdot50)}=40\ (km/h)" . Answer